A 39.1 g block of tungsten at 27.7°C and a 31.6 g block of silver at -119°C are placed together in an insulated container. (The specific heat of tungsten is 134 J/kg-K and the specific heat of silver is 236 1/kg.K.) (a) What is the equilibrium temperature? What entropy changes do (b) the tungsten, (c) the silver, and (d) the tungsten-silver system undergo in reaching the equilibrium temperature? (a) Number (b) Number (c) Number (d) Number Click if you would like to Show Work for this question: Units C Units 3/K Units J/K Units J/K Open Show Work
Answer #1

a)

mt = mass of tungsten block = 39.1 g = 0.0391 kg

ms = mass of silver block = 31.6 g = 0.0316 kg

cs = specific heat of silver = 236 J/kg-K

ct = specific heat of tungsten = 134 J/kg-K

Tti = initial temperature of tungsten block = 27.7 oC

Tsi = initial temperature of silver block = - 119 oC

Tti = initial temperature of tungsten block = 27.7 oC

T = equilibrium temperature

Using conservation of heat

Heat lost by tungsten = Heat gained by silver

mt ct (Tti - T) = ms cs (T - Tsi )

(0.0391) (134) (27.7 - T) = (0.0316) (236) (T - (- 119))

T = - 58.5 oC

or T = - 58.5 + 273 = 214.5 K

b)

St = entropy change for tungsten = mt ct ln(T/Tti) = (0.0391) (134) ln(214.5/(27.7 + 273)) = - 1.77

c)

Ss = entropy change for silver = ms cs ln(T/Tsi) = (0.0316) (236) ln(214.5/(- 119 + 273)) = 2.47

d)

S = St + Ss = - 1.77 + 2.47 = 0.7